# A Closer Look into Mixture-of-Experts in Large Language Models

### v1

@@ Appendix B Averaging Expert Neurons
Appendix C Gate Embedding
@@
Figure 8: Similarity heat maps of gate embedding (leftmost graph of each layer) along with the neuron-level similarity heat maps using averaging method. The tick numbers refer to expert indices.
@@
Since each row vector of the gate embedding determines the gate decision, there may exist some relationship between the embedding matrix of gate and of experts. To investigate this, we measure the similarities between the gate embedding vectors and compare with the averaged heat maps in Sec 4.1.2. The qualitative analyses for the combined graphs shown in Fig. 8 are detailed in this section. The table containing the $R$ values of each layer (Tab 4) is also appended at last.
Mixtral. Focusing on the heat maps of gate embedding, the similarities typically range from 0.2 to 0.4, while the values in the last layer have a rather obvious gain. Moreover, dark cross is rarely found. Surprisingly, the patterns in the heat maps of gate embeddings and of expert neurons in $W_{\text{gate}}$ are partially alike in some of the layers. This implies that the way a gate selects experts might be relevant to the way an expert activates its neurons.
DeepSeek. Unlike the almost all-zero heat maps of $W_{\text{up}}$ and $W_{\text{down}}$, the similarities of gate neurons sometimes get higher than 0.4. In addition, the heat maps of gate embeddings and of $W_{\text{gate}}$ show similar patterns. However, the overall similarities of gate embeddings decrease with depth whereas the similarities of $W_{\text{gate}}$ gradually grow. It means as layer goes deep, the gate “looks” at the input feature in more different ways when assigning scores to different experts, and meanwhile, the neuron activation of experts gets more alike.
Gork. Both dark and bright crosses commonly exist in the heat maps of gate embedding, whose patterns are similar to those of $W_{\text{gate}}$. Specially, their patterns show opposite color tendency (i.e., the deep color positions in one heat map becomes light color in another) starting form the intermediate layers. The similarities of gate embedding get lower when the layer goes deep, except for the last few layers.
Appendix D Norms of Expert Outputs and Gate Scores
@@
Figure 9: Counts of the gate score ranking for each norm ranking. The larger the rank number, the larger the norm or score.
@@
In Sec 5.2, we notice that in some MoE models, the two experts chosen by the gate usually output feature vectors with the highest norms. Therefore, we repeat the experiment using the long input and the statistical results are shown in Fig. 9.
@@
Mixtral. It can be clearly observed that the expert which outputs the largest norm is assigned the highest score most often. Surprisingly, for every $i$, the $i$-th highest score is the most likely to be assigned to the expert with the $i$-th highest output.
DeepSeek. For the experts that output the first few largest norms (rank 60th to 64th), they are most likely to be assigned the highest scores. But we do not see similar relationship for the rest of the experts. On the contrary, the gate assigns relatively high scores more frequently than low scores to the experts with the smallest norms. For the experts whose output norms rank 49th to 59th, they are commonly assigned either low scores or high scores.
@@
Grok. In contrast to the previous models, the output norms of the Grok experts tend to oppositely relate to the scores. More generally, the experts with the first few highest outputs are frequently assigned either low scores or high scores. One of the possible reasons might be the relatively low activation ratios of GeLU (see Sec 5.3) result in weaker dependence on the norm for the gate decisions.
Appendix E Chosen Experts
@@
Figure 10: Routing pattern of different models. Deeper colors mean higher gate scores assigned to the corresponding experts. Only the scores of the top-$k$ experts are illustrated.
@@
This experiment aims to examine the routing pattern. We feed an input prompt with about 64 tokens into the MoE models, and simply record the gate scores (softmax applied) of selected experts for each token. In addition to the base model of Mixtral (Mixtral-Base), we include its instruct version (Mixtral-Instruct) in this experiment. The results are depicted in Fig. 10.
@@
Mixtral. The experts in Mixtral-Base are evenly selected across tokens and it is quite common to see a sequence with more than 4 tokens being routed to a same expert. But the “special expert” with the dark cross in previous similarity graphs turns out to be an exception. These special experts are less frequently chosen and are assigned relatively low scores. The routing pattern of Mixtral-Instruct is mostly identical to Mixtral-Base, which indicates fine-tuning has little impact on gate decisions.
@@
DeepSeek. Some of the layers exist an expert that is selected by most of the tokens. However, no distinction is observed for these experts in previous similarity heat maps. Note that the gate scores of DeepSeek are typically lower than the scores of Mixtral because DeepSeek applies softmax before the top-k operation while Mixtral adopts the reverse way.
@@
Grok. The expert selection is rather even and some relatively high scores exist in the deep (>30th) layers. Same as DeepSeek, the softmax is applied before the top-k operation for Grok.

### v2

@@ Appendix D Averaging Expert Neurons
Appendix E Gate Embedding
@@
Figure 8: Similarity heat maps of gate embedding (leftmost graph of each layer) along with the neuron-level similarity heat maps using averaging method. The tick numbers refer to the expert indices.
@@
Since each row vector of the gate embedding determines the gate decision, there may be a relationship between the gate embedding matrix and that of the experts. To investigate this, we measure the similarities between the gate embedding vectors and compare them with the average heat maps presented in § 4.1.2. The qualitative analysis of the combined graphs shown in Fig. 8 is detailed in this section. The table containing the $R$ values for each layer (Tab. 5) is appended at the end.
Mixtral. Focusing on the heat maps of the gate embedding, the similarities typically range from 0.2 to 0.4, with a noticeable increase in the last layer. Moreover, dark crosses are rarely found. Surprisingly, the patterns in the heat maps of gate embeddings and expert neurons in $W_{\text{gate}}$ are partially alike in some layers. This implies that the way a gate selects experts might be relevant to how an expert activates its neurons.
DeepSeek. Unlike the almost all-zero heat maps of $W_{\text{up}}$ and $W_{\text{down}}$, the similarities of gate neurons sometimes exceed 0.4. In addition, the heat maps of gate embeddings and $W_{\text{gate}}$ show similar patterns. However, the overall similarities of gate embeddings decrease with depth while the similarities of $W_{\text{gate}}$ gradually grow. This indicates that as the layer depth increases, the gate “looks” at the input feature in more diverse ways when assigning scores to different experts, even as the neuron activations of the experts become more similar.
Gork. Both dark and bright crosses commonly exist in the heat maps of the gate embedding, whose patterns are similar to those of $W_{\text{gate}}$. Specially, their patterns show opposite color tendency (i.e., deep color positions in one heat map becomes light color in another) starting form the intermediate layers. The similarities of the gate embedding decrease when the layer depth increases, except for the last few layers.
Appendix F Additional Datasets
To ensure the universality of our findings, we repeat the experiments that require the long input (§ 5.1, § 5.2) using additional datasets. Specifically, we utilize the entire test set of WikiText-103 Merity et al. (2016) (266K tokens) and of a math dataset GSM8K Cobbe et al. (2021) (84K tokens), and 1000 sequences from the code dataset Magicoder-Evol-Instruct-110K Wei et al. (2024b) (188K tokens). As shown in Fig. 91113, the new figures of Mixtral and DeepSeek align with the previous results illustrated in the main context, even when using datasets of specific subjects like math and code (we did not test on the Grok model due to limited computation resources). These supplementary results demonstrate that our findings are general and not limited to the initial input sources.
Refer to caption Refer to caption Refer to caption Refer to caption    Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption    Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Average similarity heat maps of expert output features using (1) the long sequence, (2) WikiText-103, (3) GSM8K, and (4) Magicoder-Evol-Instruct-110K. The tick numbers refer to the expert indices.
Appendix G Norms of Expert Outputs and Gate Scores
@@
Figure 10: Counts of the gate score ranking for each norm ranking using the long input. The larger the rank number, the larger the norm or score.
Figure 11: Counts of the gate score ranking for Mixtral expert ourput norm rankings using additional datasets, namely WikiText-103 (top), GSM8K (middle), and Magicoder-Evol-Instruct-110K (bottom). The larger the rank number, the larger the norm or score.
non-MoE layer idx PPL $\downarrow$ Bench avg $\uparrow$ HellaSwag MMLU GSM8K CEval
none 8.51 39.84 61.75 41.14 13.42 43.04
1 8.50 39.42 61.88 40.55 11.83 43.41
6 8.49 38.55 62.17 39.28 12.96 39.78
12 8.51 38.21 61.46 40.24 10.54 40.61
18 8.59 37.99 61.73 39.50 11.68 39.03
24 8.58 38.94 61.33 40.45 12.05 41.91
Table 6: Model performance on various benchmarks for the dynamic expert numbers experiment. “Bench avg” refers to the average performance over the four evaluated benchmarks.
Figure 13: Counts of the gate score ranking for DeepSeek expert ourput norm rankings using additional datasets, namely WikiText-103 (top), GSM8K (middle), and Magicoder-Evol-Instruct-110K (bottom) The larger the rank number, the larger the norm or score.
@@
In § 5.2, we notice that in some MoE models, the two experts chosen by the gate usually produce feature vectors with the highest norms. To further investigate this, we repeat the experiment using the long input and additional datasets, and the statistical results are shown in Fig. 101113.
@@
Mixtral. It is evident that the expert which outputs the largest norm is most frequently assigned the highest score. Surprisingly, for every $i$, the $i$-th highest score is most likely assigned to the expert with the $i$-th highest output.
[…]
@@
Grok. In contrast to the previous models, the output norms of the Grok experts tend to have an inverse relationship with the scores. More generally, the experts with the first few highest outputs are frequently assigned either low scores or high scores. One possible explanation could be the relatively low activation ratios of GeLU (see Append H), which may result in a weaker dependence on the norm for gate decisions.
num_layers 24
vocab_size 151936
hidden_size 1024
head_dim 64
q_head 16
kv_head 4
moe_hidden_dim 640
num_shared_expert 4
num_routed_expert 64
topk 4
Table 7: Model architecture for the dynamic expert numbers experiment.
Appendix H Intermediate States of Experts
While § 5.1 focused on the final outputs of experts, we continue our analysis here by examining their intermediate outputs to examine the inner states of the experts. Given an input $x$, the intermediate state of an expert refers to the output of $\operatorname{act}(W_{\text{gate}}x)$, which is a $d_{\text{hid}}$-dimensional vector. These intermediate vectors control the activation of neurons, so we simply record them for analysis with the short input used. Mixtral, Mistral, and DeepSeek utilize SiLU as the activation function, while Grok adopts GeLU. Fig. 12 depicts the magnitude of the vectors for Mixtral across three tokens.
Common. Each figure contains some horizontal lines, indicating the presence of an “outlier” expert with either the highest or lowest activation values. Nonetheless, there is no clear relationship between these phenomena and the gate decisions.
Mixtral and Mistral. For a single token, we found that, on average, the absolute activation value of 99.6% elements in each expert exceeds 0.001 after applying the SiLU activation function. This high ratio indicates that the vast majority of neurons in an expert are activated. In Fig. 12, some vertical lines across all experts are commonly found, meaning that the $W_{\text{gate}}$ matrices of different experts assign similar activation values to neurons with the same indices. In addition, the magnitude of the intermediate states grows along with layer depth, which aligns with the observation in § 5.1.
DeepSeek. On average, 99.7% of the neurons in each expert have an absolute activation value exceeding 0.001 after applying SiLU. Vertical lines rarely exist in the DeepSeek model. Similarly, the elements in the intermediate state vectors get larger as the layer goes deeper.
Grok. With GeLU as the activation function, only 25.3% neurons per Grok expert attain an absolute activation value greater than 0.001. The activation values are generally smaller than those in Mixtral and DeepSeek. Li et al. (2022); Song et al. (2024a) suggest such difference largely stems from the distinct activation functions used. Interestingly, Song et al. (2024b) further utilize the sparsity in experts within SMoE to achieve SOTA performance when activating the same number of parameters.
Appendix I Chosen Experts
@@
Figure 14: Routing patterns of different models. Deeper colors mean higher gate scores assigned to the corresponding experts. Only scores of the top $k$ experts are illustrated.
@@
This experiment aims to examine the routing patterns. We feed an input prompt with about 64 tokens into the MoE models and record the gate scores (after applying softmax) for the selected experts for each token. In addition to the base model of Mixtral (Mixtral-Base), we also include its instruct version (Mixtral-Instruct) in this experiment. The results are depicted in Fig. 14.
@@
Mixtral. In Mixtral-Base, the experts are selected fairly evenly across tokens, and it is common to see sequences of more than four tokens routed to the same expert. But the “special expert” with the dark cross in previous similarity graphs turns out to be an exception. These special experts are chosen less frequently and tend to receive relatively low scores. The routing pattern of Mixtral-Instruct is largely identical to that of Mixtral-Base, which indicates fine-tuning has little impact on gate decisions.
@@
DeepSeek. In some layers, there is an expert selected by most tokens. However, no distinct characteristics for these experts are observed in the previous similarity heat maps. Note that the gate scores for DeepSeek are typically lower than those for Mixtral because DeepSeek applies softmax before the top-k operation, while Mixtral adopts the reverse way.
@@
Grok. The expert selection is rather even and some relatively high scores exist in ${}^{\text{th}}$) layers. Same as DeepSeek, softmax is applied before the top-k operation for Grok.